Grade 10General Physics

The resultant of two forces P and Q is equal to Q√3 and it makes an angle of π/6 with the direction of P. How do you show that P=Q or 2Q by parallelogram law

Navdeep
8 Years agoGrade 10

1 Answer

Harsha Vardhan Sai
8 Years ago
DEAR NAVDEEP,THIS ALSO INVOLVES A BIT  OF MATHEMATICIS SO MAKE A NOTE OF THIS
                                 FIRST DRAW A PARELELLOGRAM AND DRAW ITS DIAGONAL AND LABEL THEM.BY PARALELLOGRAM LAW OF VECTOR ADDITION R2=P2+Q2+2PQCOSA.
BUT GIVEN R=Q(3)1/2 BY SAS  PROPERTY Q2=P2-R2-2PRCOSA
ON SUBSTITUTING VALUES Q2=P2-3Q2-2PR(3)1/2*(3)1/2/2
P2+2Q2-3PQ=0
BY FACTORISATION
(P-Q)(P-2Q)=0
THEN P=Q OR P=2Q
HENCE PROVED